OpenStudy (anonymous):

What is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 48 ?

OpenStudy (anonymous):

128: 8, 8+5, 8+ 2(5), ... 8+ (n-1) 5; n=25, n-1= 24; 8 +24(5)= 8 + 120 = 128; I have forgotten the formula; I'd have to work it out.

OpenStudy (anonymous):

THANKS!!!